# Traffic Flow and Speed Monitoring Based On Optical Fiber Distributed Acoustic Sensor

### v1

@@ IV-B Metrics
To make statistics on the characteristics of traffic flow, it is necessary to study the three fundamental indicators: velocity (V), flow (Q), and density (K) on road profile and road segment respectively.
@@
On the road profile, traffic flow $\rm Q$ is defined as the number of vehicles N passing through a given profile of the road in a given period T: $\rm{Q=\frac{N}{T}}$. The average velocity of the vehicles passing through a given point is called Time Mean Speed (TMS). According to space-time accumulation, TMS on the profile should be the harmonic average velocity, which is a spatial concept. Therefore, TMS refers to the harmonic average of the location velocity of all vehicles passing through a profile of the road in a certain observation time: $\overline{\rm V}_{\rm{TMS}}=\frac{1}{\frac{1}{N}\sum_{i=1}^{N}\frac{1}{\rm V_{i}}}$. Density is expressed as occupancy rate, i.e., time density of vehicles, which represents the ratio of the time occupied by vehicles passing a certain road profile to the observation time in a certain observation time: $\rm K=\frac{\rm Q}{\overline{V}_{\rm{TMS}}}$.
@@
On the road segment, density refers to the spatial density set of vehicles, that is, the number of vehicles N existing per unit road length L at a certain moment is $\rm K=\frac{\rm N}{\rm L}$. The average velocity of the vehicles in a given section is known as Space Mean Speed (SMS). According to space-time accumulation, SMS on the road segment should be the arithmetic average, which is a concept in time. Therefore, SMS refers to the arithmetic average of all vehicle instantaneous velocities in a certain observation section: $\overline{\rm V}_{\rm{SMS}}=\frac{\sum_{i=1}^{N}\rm V_{i}}{N}$. SMS is always lesser than or equal to TMS. Traffic flow refers to the ratio of the total distance of all vehicles to the accumulation of space and time in a certain period, $\rm Q=\rm K\overline{\rm V}_{\rm{SMS}}$.
IV-C Quantitative analysis
We tested the above algorithms on real highways. Our monitoring focuses on the target areas, and the algorithm proposed in this paper relies solely on optical fiber stripe data, independent of camera data. Consequently, when evaluating accuracy, our algorithm is compared only against optical fiber stripe data. By applying threshold processing, we can also distinguish between different vehicle classes, such as cars and trucks. Fig. 13 and Fig. 14 depict vehicle trajectory waterfall diagrams extracted by our algorithm with red lines indicating trucks and yellow lines indicating cars. Each line is uniquely identified with an individual ID.
Figure 13: An example of vehicle tracking results in the Jingjintang expressway dataset. (a) original waterfall diagram, (b) the red lines represent the extracted trajectory of trucks, and the yellow lines represent cars.
@@ IV-C Quantitative analysis
Figure 14: An example of vehicle tracking results in Jiurui tunnel dataset of short distance. (a) original waterfall diagram, (b) the red lines represent the extracted trajectory of trucks, and the yellow line represents a car.
Figure 15: Vehicle trajectory extraction results using the proposed method, Hough transform and Radon transform. The last two rows show the results using the Hough transform and Radon transform respectively.
Furthermore, We carried out comparative experiments with the Hough transform and Radon transform, employing the same preprocessing steps. As depicted in Fig. 15, our algorithm shows superior accuracy in extracting vehicle trajectories, even when these lines intersect or overlap, effectively distinguishing and separating individual vehicles.
We visually compared the algorithm results with optical fiber stripes in waterfall diagrams. The optical fiber dataset offers vehicle counts for both cars and trucks. Fig. 16 and Table. IV present a comparison of traffic flow data between the algorithm’s test results and the optical fiber results over two hours in the Jingjintang expressway dataset. Solid lines represent the vehicles detected by the algorithm, while dashed lines represent the optical fiber counts, which slightly deviate from the actual camera-based vehicle counts.
@@
TABLE III: Comparison of algorithm test counts and optical fiber stripe counts in 7 sets of Jiurui tunnel dataset.
@@
(a) Trucks
(b) Cars
Figure 16: Vehicle statistics counts in Jingjintang expressway dataset. The blue and orange lines show the counting results of the fiber stripe and the algorithm proposed in this paper respectively.
TABLE IV: Comparison of algorithm test counts and optical fiber stripe counts within 2 hours of Jingjintang expressway dataset.
Vehicle type Algorithm test counts Fiber stripe counts Accuracy rate
Truck 138 241 57.26%
Car 268 381 70.34%
Total 406 622 65.27%
The results from the Jiurui tunnel dataset are listed in Table. III. Among them, there were 143 instances of optical fiber stripes, and the algorithm’s detection results were 130 vehicles, yielding an accuracy of 90.90%. It is worth mentioning that the algorithm proposed in this paper not only exhibits consistency in counting but also maintains a high degree of alignment with vehicle trajectories through optical fiber stripes.
In addition to vehicle trajectory extraction, our approach can perform comprehensive statistical analyses, including velocity, flow and density on both road profiles and road segments. For example, based on the extraction results of structured vehicle data from the dataset in Fig. 14 of the tunnel highway scenario, we can specify a certain road profile or a certain segment and make traffic flow information statistics like Table. V.
TABLE V: Some traffic flow information statistics.
@@
V Conclusion
In this paper, we have demonstrated the feasibility of employing optical fiber Distributed Acoustic Sensors(DAS). It can serve as a valuable supplement to existing traffic detection technologies and offers a feasible scheme for achieving long-distance vehicle trajectory tracking and comprehensive space-time traffic detection objectives for intelligent highways. However, there are certain deficiencies and areas for improvement in this research, including:
(1) Handing heavy traffic and parallel passages: Our proposed vehicle detection method works effectively for single-file traffic scenarios. However, in cases of heavy traffic with large and small vehicles passing in parallel, improving how we distinguish vehicle counts and precisely identify vibration signal positions is necessary.
@@ V Conclusion
(3) Considering complex traffic scenarios: To deploy fiber-assisted intelligent transportation successfully, we also need to address complex vehicle behaviors in real highway conditions, such as overtaking and stopping, to provide early warnings of emergencies.
@@ Acknowledgments
This research was funded by Tianjin Yunhong Technology Development Grant No. 2021020531.
This work was supported in part by the National Natural Science Foundation of China Grant No. 61972404.

### v2

@@ 4.2.2 Traffic indices
To make statistics on the characteristics of traffic flow, we utilize three fundamental indices: velocity (V), traffic flow (Q), and density (K) on road profile and road segment respectively.
@@
On the road profile, $\rm Q$ is defined as the number of vehicles $\rm N$ passing through a given profile of the road in a given period $\rm T$: $\rm{Q=\frac{N}{T}}$. The average velocity passing through a given point is called time mean speed (TMS). TMS refers to the harmonic average of the location velocity of all vehicles passing through a profile of the road in a certain observation time: $\overline{\rm V}_{\rm{TMS}}=\frac{1}{\frac{1}{\rm N}\sum_{i=1}^{\rm N}\frac{1}{\rm V_{i}}}$. Density is expressed as occupancy rate, i.e., time density of vehicles, which represents the ratio of the time occupied by vehicles passing a certain road profile to the observation time in a certain observation time: $\rm K=\frac{\rm Q}{\overline{V}_{\rm{TMS}}}$.
@@
On the road segment, $\rm K$ is the spatial density set of vehicles, that is, the number of vehicles $\rm N$ existing per unit road length $\rm L$ at a certain moment is $\rm K=\frac{\rm N}{\rm L}$. The average velocity of the vehicles in a given section is known as space mean speed (SMS). SMS refers to the arithmetic average of all vehicle instantaneous velocities in a certain observation section: $\overline{\rm V}_{\rm{SMS}}=\frac{\sum_{i=1}^{\rm N}\rm V_{i}}{\rm N}$. Traffic flow refers to the ratio of the total distance of all vehicles to the accumulation of space and time in a certain period, $\rm Q=\rm K\overline{\rm V}_{\rm{SMS}}$.
4.3 Analysis
We tested the proposed algorithms on real tunnels and cities. In tunnels, the wavelet denoising threshold $\lambda=0.15$ and $a=0.5$ in Eq. (7). Butterworth low-pass filter was applied to the first col signals with order $N=1$ and normalized cut-off frequency $W_{n}=0.5$. Then we extracted signal peaks using the minimum height of 0.06. In the phase of vehicle tracking, the initial velocity interval is $v_{\min}^{(1)}=60$ km/h to $v_{\max}^{(1)}=120$ km/h. The vehicles with an average peak value above 0.5 and width above 3 were trucks and the rest were cars. Fig. 15 depicts ablation experiments regarding different denoising algorithms. In these vehicle trajectory waterfall diagrams, each vehicle is uniquely identified with an individual ID. We conducted quantitative comparative experiments on the effectiveness of the denoising algorithm. As shown in Table. 4, improved threshold denoising method has smaller MSE, larger PSNR and SSIM, indicating better results.
(a) Without denoising.
(b) Soft threshold denoising.
(c) Improved threshold denoising.
@@ 4.3 Analysis
Figure 15: An example of vehicle tracking results in Jiurui tunnel dataset of short distance. The red lines represent the extracted trajectory of trucks while the yellow line represents a car. (a) Trajectory without denoising, (b) trajectory using soft threshold denoising, and (c) trajectory using improved threshold denoising.
Table 4: Quantitative indexes of different denoising algorithms.
Method MSE PSNR SSIM
Soft wavelet 0.0568 73.0321 0.8645
Improved wavelet 0.0356 77.0969 0.9161
Table. 5 present results of algorithm test counts and fiber stripe counts in Jiurui tunnel dataset. Among them, there were 143 instances of fiber stripes, and the algorithm’s detection results yielded an accuracy of 90.90%. Even on certain road sections, the accuracy rate can reach 100%.
@@
Table 5: Comparison of algorithm test counts and optical fiber stripe counts in 7 sets of Jiurui tunnel dataset.
@@
Furthermore, after employing the same preprocessing steps, we carried out comparative experiments with Hough transform [13, 14] and Radon transform [15] under complicated events like congestion and overtaking events. As depicted in Fig. 16, Hough transform identifies a single vehicle trajectory line as multiple lines; Radon transform overlooks detecting lines with short driving distances. Our algorithms can effectively distinguish and separate individual vehicles, showing superior consistency and accuracy even when lines intersect or overlap.
(a) Hough transform.
(b) Radon transform.
(c) Our proposed method.
Figure 16: Vehicle trajectory extraction results. (a) Hough transform, (b) Radon transform, and (c) our proposed method.
Considering that our experiments are only carried out on highways, we conduct comparative experiments on public datasets of normal urban traffic in [12]. Besides, we compare our proposed algorithm with deep deconvolution and MUSIC beamforming algorithms[12] to further prove the effectiveness of our proposed algorithms. As shown in Fig. 17, we selected three different representative fiber signals in datasets of [12] for experiments. It is worth mentioning that our algorithm can better extract the entry time and velocity information of vehicles, and can track different vehicles throughout the whole process.
Figure 17: Three examples of different traffic conditions, ordered by increasing difficulty. The top panels (a, b, c) show the original DAS recordings, the panels (d, e, f) and (g, h, i) show the results of deep deconvolution and proposed algorithms respectively. Furthermore, our algorithm can track the whole trajectories, as shown in panels (j, k, i).
In addition to vehicle trajectory extraction, our approach can perform comprehensive statistical analyses, including velocity, flow and density on both road profiles and road segments. For example, based on the extraction results of structured vehicle data from the dataset in Fig. 15 of the tunnel scenario, we can specify a certain road profile or a certain segment and make traffic flow information statistics like Table. 6.
Table 6: Some traffic flow information statistics.
@@
5 Perspectives
Despite the advantages of DAS systems, such as low cost, precise position, and corrosion resistance, it does have inherent limitations. As explained in Section 4.1.1, there are several complex traffic situations difficult to deal with, including traffic flow with cars and trucks, multi-vehicle parallel, and opposite interference. Having described the potential of DAS systems to promote ITS, we further explore the possibility of DAS systems and camera systems integration to achieve better results.
The video-based multi-object tracking (MOT) algorithm is divided into two stages: object detection and object tracking. In the object detection stage, we utilized deep learning-based YOLOv8 [38] algorithm and fine-tuned it on the UA-DETRAC dataset [39]. In the object tracking stage, we applied the BoT-SORT algorithm [40] for subsequent object tracking, which is a multi-object tracking algorithm that uses a two-stage Kalman filter and object feature matching. By integrating the YOLOv8 and BoT-SORT algorithms, we can obtain vehicle tracking results from video-based systems. Further, we transform the video perspective projection to the aerial perspective to compare with fiber-based systems using perspective transformation [41]. The perspective matrix here is:
$\begin{pmatrix}-0.051&-0.007&75.813\\ 0&-0.041&44.298\\ 0&0.007&1\end{pmatrix}$ (17)
As shown in Fig. 18, two vehicles passed through within 30 seconds with a tanker in front and a car behind. The two red lines represent the fiber-based tracking results of the two vehicles. Due to the length of the tanker, only a few frames can be detected, making it impossible to track the entire trajectory in vision-based system. Only the car can be tracked throughout the process. After perspective transformation, the car’s trajectory is projected onto the optical fiber diagram, as shown by the black line, with bidirectional arrows indicating the corresponding spatial relationship between the video and the fiber images. It is worth mentioning that due to the continuous laying of optical fiber, vehicles can be continuously tracked, unrestricted by the fixed monitoring angle of cameras. Since DAS systems can only provide accurate vehicle location information rather vehicle and road conditions visualization, we can consider providing more accurate information through the video perspective, such as lane, model, license plate, color and other information. Therefore, in future work, the combination of video-based and fiber-based systems can be furthur explored to promote the further development of ITS.
Figure 18: Fiber-based and vision-based tracking results. In the waterfall diagram in the middle, the two red lines represent the DAS tracking results of the two vehicles. A black line represents the video tracking track of the following line of vehicles after video detection and tracking, and describes the corresponding relationship between the positions of vehicles in the three pictures.
6 Conclusion
In this paper, we have demonstrated the feasibility of employing DAS systems for traffic monitoring. Considering the challenges when facing vehicle complex behaviors like congestion and illegal overtaking, we designed a series of signal processing and feature extraction to achieve multi-vehicle tracking. The data preprocessing utilizes a wavelet threshold denoising algorithm combining hard and soft threshold functions. Based on preprocessed optical fiber waterfall diagrams, we propose a vehicle position detection algorithm using the Butterworth low-pass filter and peaks location search. Standard indices MSE, PSNR and SSIM were employed for quantitative analysis to enhance data comprehension. Then a line-by-line match algorithm is proposed for vehicle trajectory extraction to achieve overall multi-vehicle tracking. Using the extracted vehicle information, the velocity, traffic flow and density of any specified road profile and road segment can be further calculated. Experiments were carried out in highway, tunnel and city scenarios. In the simple section of tunnel scenario, the algorithm’s detection results yielded an accuracy of 90.90%. In some complex sections, compared with the classical Hough and Radon transform-based methods and MUSIC beamforming algorithm, the visualized results of our proposed methods demonstrate advantages in tracking vehicles and measuring speed in long-distance.
However, there are certain deficiencies and areas for improvement in this research, including:
(1) Handing heavy traffic and parallel passages: Our proposed vehicle detection method works effectively for simple traffic scenarios. However, in cases of heavy traffic with large and small vehicles passing in parallel, improving how we distinguish vehicle counts and precisely identify vibration signal positions is necessary.
@@ 6 Conclusion
(3) Considering complex traffic scenarios: To deploy fiber-assisted intelligent transportation successfully, we also need to address more complex vehicle behaviors in real highway conditions, such as overtaking and stopping, to provide early warnings of emergencies.
Overall, our study demonstrates the great potential of DAS applications in traffic monitoring. In future work, we expect to integrate the advantages of multiple sensors (such as fibers, videos and cameras) and further serve intelligent transport systems.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
[…]
@@ Acknowledgments
This research was funded by Tianjin Yunhong Technology Development (Grant number: 2021020531).